OpenStudy (anonymous):

What is the inverse of the exponential function, f(x)=5^x?

OpenStudy (zzr0ck3r):

\(y=a^x\implies \log_a(y)=x\) so, we just switch the variables \(y=\log_a(x)\)

OpenStudy (zzr0ck3r):

\(f(x)=y=a^x\implies \log_a(y)=x\) so we make this our inverse \(f^{-1}(x)=\log_a(x)\)

OpenStudy (anonymous):

So it would just be log_5^x?

OpenStudy (anonymous):

hmmm?

OpenStudy (zzr0ck3r):

write it like this \(\log_5(x)=\) log_5(x)

OpenStudy (zzr0ck3r):

dont use ^

OpenStudy (anonymous):

Cool, thank you

OpenStudy (zzr0ck3r):

np